# Order of Addition in Mixture-Amount Experiments

### v1

@@ 3.2 Construction of OofA component-amount design
A three ingredients component-amount design with five levels of the total amount $(A=0,1/2,2/3,\\ 3/4,1)$ is developed by projecting a four-component simplex centroid design. This is done by following the footstep of Prescott and Draper (2004). Using this as the base design, we derived a three-components OofA component-amount design with 31 design points and included three PWO variables $z_{12}$, $z_{13}$, and $z_{23}$. By way forward, we can fit the models (7) and (8), and can analyze the design. The constructed design is given in Table 2.
Table 2: The OofA mixture component-amount design for three ingredients and five levels of total amount $A$.
Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ $A$
@@ 4.1 Study of the joint action of three related hormones in groups of mice
This example is taken from Cornell (2002). The presented results of an experiment involved administering $10$ different blends or combinations of three distinct hormones to $10$ groups of $12$ mice each. The detail can be found in Claringbold (1955). The hormone blends were also prepared in three amounts: $0.75\times 10^{-4}\mu g$, $1.50\times 10^{-4}\mu g$, and $3.00\times 10{-4}\mu g$, resulting in a total of $30$ groups. The primary focus is on determining how different mixtures of hormones influenced the cornification response of the vaginal epithelium in ovariectomized mice. As we consider the order of addition in this experiment, therefore using $\{3,3\}$ simplex-lattice design, we get 63 blends in the hormone proportions $x_{1},x_{2},x_{3}$ with PWO variables $z_{12},z_{13},z_{23}$, and with the total amounts $A=0.75,1.50,3.00$. The design matrix, using model (6), is given in Table 3. The objective of the analysis includes how the order of addition and proportions impacted the response while fitting the mixture-amount model (6) in the design matrix given in Table 3. For model identifiability, we omit the mixture-order interactions $x_{1}z_{13},x_{2}z_{12},x_{3}z_{23}$ and only use $x_{1}z_{12},x_{2}z_{23},x_{3}z_{31}$.
Table 3: The OofA simplex-lattice design with $m=3$ and $l=3$ with three levels of total amount $A$
Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ $A$ Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ $A$
@@
The design analysis reveals that the maximum prediction variance of the design is 0.761, which is recommended for moderate prediction accuracy. The average prediction variance of the design is 0.571, which suggests a reasonably good overall predictive performance that balances accuracy and robustness. A G-efficiency of 75.1% is pretty good, indicating that the design is quite efficient. This means the design is well-constructed, balancing the number of runs and the precision of parameter estimates. We use the fraction of design space plot (FDS) to assess design prediction capability visually: see Figure 1. The FDS plot indicates that the design is robust, with a large portion of design space having low prediction variance. This suggests reliable and consistent predictions for most factor combinations.
Figure 1: Fraction of Design Space plot for OofA mixture-amount model.
We performed a power analysis in Table 4 to determine the likelihood of detecting significant effects in the experiment.
Table 4: Power analysis at 5% alpha level for effects
Terms Std. Error $\frac{1}{2}$ Std. Dev Terms Std. Error $\frac{1}{2}$ Std. Dev
$x_{1}$ 0.36 10.2% $x_{1}x_{2}$ 0.57 7.1%
$x_{2}$ 0.36 10.2% $x_{1}x_{3}$ 0.57 7.1%
$x_{3}$ 0.36 10.2% $x_{2}x_{3}$ 0.57 7.1%
$z_{12}$ 0.28 13.5% $x_{1}z_{12}$ 0.84 6.0%
$z_{13}$ 0.27 14.2% $x_{2}z_{23}$ 0.73 6.3%
$z_{23}$ 0.27 14.2% $x_{3}z_{13}$ 0.73 6.3%
The standard errors for the main effects and the $1/2$ standard deviation values indicate moderate precision in estimating these effects. All three hormones and their order of addition in the mixture significantly affect the cornification response. The interaction effects between each pair of the three hormones are significant, as are the interaction effects between the hormones and the PWO variables. However, the precision of these estimates is lower than that of the main effects.
@@ 4.2 A Placebo tablet formulation with three components
It might be easier to explain dosage units like tablets or gelatin capsules by the amount of each component rather than their proportions when the total mass varies. However, using proportions and the variable total amount in the model is usually more convenient. This example is taken from Lewis et al. (1991). Huisman et al. (1984) have discussed using a mixture of experimental designs to prepare placebo tablets. Their formulation included up to three diluents: $\alpha$-lactose monohydrate, potato starch, and anhydrous $\alpha$-lactose. The tablets also had magnesium stearate as a lubricant, but since this was kept at a constant low level, the rest of the formulation, made up entirely of diluents, was considered to be 100%. The measured response was the crushing strength of the tablets. In practice, the total weight of the three components (diluents) in a placebo tablet formulation is typically up to $500mg$ i.e. $A_{max}=500mg$. We assume that several dosage strengths are to be formulated. A component-amount model (8), with interaction effect and including PWO variable, is used. With the total maximum, the design given in Table 2 is transformed this amount and the resulting design is given in Table 5. The design has five levels of the total amount, i.e., $0,250,333.3,375$, and $500$. The zero-amount situation is used for control test and discussed by Piepel (1988).
Table 5: The OofA mixture component-amount design for three ingredients and five levels of total amount $A$.
Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ $A$ Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ $A$
@@
Once we analyse the design, we notice that the maximum prediction variance is $0.96$, which ensures that most predictions remain reasonably reliable. A value of 0.516 for average prediction variance suggests that, on average, model predictions are reasonably precise. A G-efficiency of 53.8% is above the commonly accepted threshold of 50%, indicating that the design is relatively efficient. The Scaled D-optimality of $12.29$ indicates that the design is relatively efficient in estimating model parameters.
Table 6: Power at 5% alpha level for effects
Terms Std. Error 2 Std. Dev $R_{i}^{2}$ Terms Std. Error 2 Std. Dev $R_{i}^{2}$
$a_{1}$ 2.17 7.2% 0.9645 $a_{33}$ 0.96 49.8% 0.7966
$a_{2}$ 2.17 7.2% 0.9645 $a_{1}a_{2}$ 1.48 9.7% 0.9127
$a_{3}$ 2.17 7.2% 0.9645 $a_{1}a_{3}$ 1.48 9.7% 0.9127
$z_{12}$ 0.63 32.0% 0.8416 $a_{2}a_{3}$ 1.48 9.7% 0.9127
$z_{13}$ 0.63 32.0% 0.8416 $a_{1}z_{12}$ 1.69 8.6% 0.8527
$z_{23}$ 0.63 32.0% 0.8416 $a_{2}z_{23}$ 1.69 8.6% 0.8527
$a_{11}$ 0.96 49.8% 0.7966 $a_{3}z_{13}$ 1.69 8.6% 0.8527
$a_{22}$ 0.96 49.8% 0.7966
The standard errors for linear effects $a_{1},a_{2}$ and $a_{3}$ are relatively high, suggesting that all three diluent effects are less significant. High $R^{2}$-value $(0.9645)$ suggests that these effects explain a significant portion of the variability in the response. The quadratic effects $(a_{11},a_{22},a_{33})$ and certain interactions $(a_{1}a_{2},a_{1}a_{3}$ and $a_{2}a_{3})$ show possible significance. The orders of addition effects $(z_{12},z_{13}$ and $z_{23})$ are likely to be significant, especially at larger standard deviations. The interaction effects between diluents and their order of addition $(a_{1}z_{12},a_{3}z_{13}$, and $a_{2}z_{23})$ may be less significant but still warrant consideration.
The FDS plot in Figure 2 shows that the majority of the design space has a prediction variance below 0.96. The higher prediction variances are limited to a smaller portion of the design space. So FDS plot for this design likely indicates strong prediction capabilities across most of the design space.
Figure 2: Fraction of Design Space Plot for OofA component-amount model.
@@ 5 Conclusion
In this study, we developed OofA mixture-amount experiments where the component amounts and the order of their addition to the mixture affect the overall response. We modified mixture-amount and component-amount models by incorporating the order-of-addition effect. To fit these models, we constructed designs for the OofA mixture-amount experiment. The designs constructed in this paper can be used to determine whether the order and the mixture amounts have a significant effect on the response. These full mixture-amount OofA designs confirmed orthogonality between mixture-amount model terms and OofA effects. This allows precise estimation of both mixture-component parameters and order-of-addition effects.
In the first example, we constructed a three-ingredient OofA mixture-amount design using three levels of mixture total A. We fitted the model (6) and analyzed that the model has $35$ and residuals have $27$ degrees of freedom, which allows for estimating main effects, interactions, and higher-order interactions and a reasonable basis for error estimation. The G-efficiency value $75.1\%$ indicates that the design is quite efficient. Further, it was noted that proportions and order-of-addition effects are significant. The interactions of mixture components and order-of-addition effects exhibit higher variability and potentially lower significance.
In the second example, we generated a three-components OofA component-amount design in 31 runs with five levels of total amount. We fitted the OofA component-amount model given in (8) to the design derived. The analysis described that, on average, the model predictions are reasonably precise, fairly robust in parameter estimation, and rational level of variability in parameter estimates. A G-efficiency of 53.8% indicates that the design is relatively efficient. The component amounts, order of addition, and their interaction effects are significant.
The developed OofA designs provide a systematic approach to including order-of-addition effects in mixture-amount experiments, enabling a more comprehensive design space exploration and better optimization of the response variable. It is the initial step for designing OofA mixture-amount experiments. For future work, it would be best to use a fraction of such designs to control the cost of the experiment. This work can be extended to the orthogonally blocked mixture experiments.
Acknowledgment
Touqeer Ahmad acknowledges support from the Règion Bretagne through project SAD-2021- MaEVa.

### v2

@@ 3.2 Construction of OofA component-amount design
A three-instrument component-amount design with five levels of the total amount $(A=0,1/2,2/3,3/4,1)$ is developed by projecting a simplex centroid design with four ingredients. This is done by following the methods used in [16]. Using this as the base design, we derived a three-component OofA component-amount design with 31 design points and included three PWO variables $z_{12}$, $z_{13}$, and $z_{23}$.The constructed design is given in Table 3. In addition, we can fit the models (6) and (7), and we can analyze the design.
Table 3: The OofA mixture component-amount design for three ingredients and five levels of total amount AA.
Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA
@@
A general construction of OofA mixture-amount and component amount design is given in the Algorithm 3.
Algorithm 3 General Construction of OofA Mixture-Amount and Component-Amount Design Matrices
1:Input: Number of ingredients $m$, total amount set AA, initial design matrix $D$.
2:Output: Final design $D^{*}$ with improved G-efficiency.
3:Let $\mathcal{C}$ be the set of all $m!$ permutations of ($c_{1},\dots,c_{m})$.
4:For each $1\leq j<k\leq m$, define pairwise order variables:
[…]
5:Construct extended design space:
$\mathcal{D}_{\text{OofA}}=\left\{(\mathbf{x},a,\mathbf{c},\{z_{jk}\}):\mathbf{x}\in\Delta^{m-1},a\in A,\mathbf{c}\in\mathcal{C}\right\}$
6:Apply TA-algorithm 2 to find G-optimal design with reduced runs.
7:Output: Optimized OofA design matrix $D^{*}$.
@@ 4.1 Study of the joint action of three related hormones in groups of mice
This example is taken from Cornell [2]. The presented results of an experiment involved administering $10$ different blends of three distinct hormones to $10$ groups of $12$ mice each. The detail can be found in Claringbold [24]. The hormone blends were also prepared in three amounts: $0.75\times 10^{-4}\mu g$, $1.50\times 10^{-4}\mu g$, and $3.00\times 10{-4}\mu g$, resulting in a total of $30$ groups. The primary focus is to determine how different mixtures of hormones influenced the cornification response of the vaginal epithelium in ovariectomized mice. The response ($y$) referred to the percent response, which represents the proportion of mice (out of 12) in each group that responded to hormone injection. It was measured for each combination of mixtures at different levels of total hormone amount AA $(0.75,1.5$ and $3.0)$. The $10$ hormone blends at each amount $A$ form a $\{3,3\}$ simplex lattice, creating a $\{3,3\}$ x $3$ mixture amount design. A quadratic mixture-amount model  (1) was fitted and it was observed that the effect on the percent response of increasing the amount of hormone $3$ is less than with hormone $1$ and hormone $2$. Now we consider the order of addition in the experiment. So the OofA $\{3,3\}$ simplex-lattice design now has $21$ mixture blends ($x_{1},x_{2},x_{3}$$z_{12},z_{13}$ and $z_{23}$) including PWO variables at each total amount $A$.This creates a $63$ blend OofA mixture-amount design given in Table 4. The objective of the analysis includes how the proportions and order of addition impacted the response when we fit the mixture-amount model (5). For model identifiability, the mixture-order interactions $x_{1}z_{13},x_{2}z_{12},x_{3}z_{23}$ are ignored.
Table 4: The OofA simplex-lattice design with $m=3$ and $l=3$ with three levels of total amount AA
Runs $x_{1}$ $x_{2}$ $x_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA Runs $x_{1}$ $x_{2}$ $x_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA
@@
Since the design has two many runs, we select a subset of runs from our full design. We implemented the TA algorithm in Python to generate a G-optimal OofA mixture-amount design using 10,000 iterations. The G-optimal OofA mixture amount design generated via the TA algorithm. The design has 36 runs, which is the minimum runs required to estimate the model parameters  (5). We simulated the percent responses ($y$) for each mixture based on the results of Claringbold [24], Cornell [2].The design is given in Table 5. The simulated responses are consistent with the findings reported in Claringbold [24], Cornell [2]. We fit the OofA mixture-amount model (5) to the design given in Table 5. We used ANOVA to assess the significance of the effects of mixture proportions with variable total amounts and their order-of-addition effect on the response. The results are presented in Table  6. At the 0.05 significance level, it can be observed that both the hormone proportions at different levels of total amount and their order of addition have a significant effect on the percentage response (proportion of mice out of 12). The analysis of the design reveals that the maximum prediction variance of the design was found to be 1.00, while the average prediction variance of the design was 0.89. A G-efficiency of 89% implies that the design performs well in terms of minimizing the worst-case prediction variance relative to the average prediction variance. The design provides reliable and stable predictions throughout the mixture space, making it a strong candidate to accurately model response surfaces.
Table 5: G-optimal OofA mixture component–amount design obtained via TA algorithm
Run $x_{1}$ $x_{2}$ $x_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA $y$ Run $x_{1}$ $x_{2}$ $x_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA $y$
1 0 0.67 0.33 0 0 1 3.0 56 19 1 0 0 0 0 0 0.75 40
2 0 0 1 0 0 0 1.5 52 20 0.67 0.33 0 -1 0 0 3.0 15
3 0.67 0.33 0 1 0 0 0.75 29 21 0.67 0 0.33 0 1 0 3.0 54
4 0.67 0.33 0 -1 0 0 0.75 95 22 0 0.33 0.67 0 0 1 3.0 19
5 1 0 0 0 0 0 3.0 19 23 0 1 0 0 0 0 0.75 59
6 0.33 0 0.67 0 1 0 0.75 38 24 0.33 0.33 0.33 1 -1 -1 3.0 25
7 0.33 0 0.67 -1 0 0 0.75 26 25 0.67 0.33 0 1 0 0 0.75 94
8 0 1 0 0 0 0 0.75 50 26 0.33 0.33 0.33 1 1 1 0.75 42
9 0 0.33 0.67 0 0 1 0.75 50 27 0.33 0.33 0.33 -1 1 1 3.0 76
10 0.33 0 0.67 0 -1 0 3.0 63 28 0 1 0 0 0 0 1.5 34
11 0 1 0 0 0 0 0.75 95 29 0.33 0.33 0.33 1 1 1 3.0 45
12 0.33 0 0.67 0 1 0 1.5 45 30 0 1 0 0 0 0 3.0 43
13 0.67 0.33 0 -1 0 0 1.5 33 31 0.33 0.33 0.33 -1 -1 1 3.0 60
14 0 0.33 0.67 0 0 -1 0.75 22 32 0.33 0.33 0.33 1 -1 -1 3.0 43
15 0 0.33 0.67 0 0 -1 1.5 51 33 0.67 0 0.33 0 1 0 1.5 42
16 0.67 0.33 0 1 0 0 1.5 39 34 0.33 0.33 0.33 -1 1 1 1.5 35
17 0.33 0.33 0.33 -1 -1 -1 1.5 36 35 0.33 0 0.67 0 1 0 3.0 40
18 0 0 1 0 0 0 1.5 73 36 0.33 1 0 0 0 0 3.0 9
Table 6: Analysis and parameter estimates of OofA mixture–amount model
Term Estimate Std.Err t value Term Estimate Std.Err t value
$x_{1}$ -73.28 10.56 -6.940 $z_{13}A$ -6.32 43.84 -0.144
$x_{2}$ -45.92 17.15 -2.667 $z_{23}A$ -30.19 43.54 -0.693
$x_{3}$ -35.78 13.06 -2.737 $x_{1}x_{2}A$ -4.11 26.82 -0.153
$z_{12}$ 31.22 14.52 2.388 $x_{1}x_{3}A$ 194.43 171.28 1.135
$z_{13}$ -7.13 24.19 -0.294 $x_{2}x_{3}A$ 56.77 44.10 1.287
$z_{23}$ -24.14 11.98 -2.015 $x_{1}z_{12}A$ -5.68 42.15 -0.135
$x_{1}x_{2}$ -45.21 37.21 -1.215 $x_{2}z_{23}A$ 21.31 20.28 1.050
$x_{1}x_{3}$ 47.77 30.05 2.589 $x_{1}A^{2}$ -15.73 12.64 -1.224
$x_{2}x_{3}$ -45.66 57.60 -0.792 $x_{2}A^{2}$ -48.82 83.10 -0.527
$x_{1}z_{12}$ -25.71 11.41 -2.253 $x_{3}A^{2}$ 773.37 719.76 1.070
$x_{3}z_{13}$ -8.96 69.19 -0.129 $z_{12}A^{2}$ 151.87 56.28 2.690
$x_{2}z_{23}$ -23.05 14.18 -1.625 $z_{13}A^{2}$ -5.59 42.37 -0.132
$x_{1}A$ -14.85 11.79 -1.259 $z_{23}A^{2}$ -33.46 41.37 -0.808
$x_{2}A$ 53.47 28.64 1.866 $x_{1}x_{2}A^{2}$ -19.85 62.96 -0.315
$x_{3}A$ 753.39 768.98 0.979 $x_{1}x_{3}A^{2}$ -221.50 177.09 -1.250
$z_{12}A$ 128.99 64.24 2.008 $x_{2}x_{3}A^{2}$ 68.41 77.64 0.881
@@
For visual assessment of the prediction capability of the design, we use the fraction of the design space plot (FDS), as shown in Figure 1. The FDS plot indicates that the design is robust, with a large portion of design space having low prediction variance. This suggests reliable and consistent predictions for most factor combinations.
Figure 1: FDS Plot for OofA mixture-amount model
A simulation study using 10,000 synthetically generated datasets demonstrated that the reduced 36-run design, obtained by the Threshold Accepting (TA) algorithm, maintained predictive accuracy comparable to the full 63-run design. Although the G-efficiency of the full design was higher (4520.3) due to redundancy, the TA design consistently exhibited strong mean G-efficiency (2451.3), with a substantially reduced experimental burden (36 design points). The mean normalized G-efficiency for TA Design is 51.3% which means that the 36-point design still retains about 51% of the prediction capability of the full design. This validates the practical utility of the TA-algorithm based design, showing that it can deliver robust performance with fewer experimental runs.
Figure 2: Boxplot of MSEs across 10,000 simulated datasets
The boxplot of MSEs for 10,000 simulations in Figure  2 shows that the TA-reduced design (36 points) achieves a lower median MSE compared to the full design (63 points), while exhibiting a slightly wider spread of errors. This indicates that the TA algorithm can efficiently select design points, maintaining predictive accuracy with substantially fewer runs.
@@ 4.2 A Placebo tablet formulation with three components
It may be easier to explain dosage units such as tablets or gelatin capsules by the amount of each component rather than their proportions when the total mass varies. However, using proportions and the variable total amount in the model is usually more convenient. This example is taken from Lewis et al. [25]. Huisman et al. [26] have discussed the use of a mixture of experimental designs to prepare placebo tablets. Their formulation included up to three diluents: $\alpha$-lactose monohydrate, potato starch, and anhydrous $\alpha$-lactose. The tablets also had magnesium stearate as a lubricant, but since this was kept at a constant low level, the rest of the formulation, made up entirely of diluents, was considered 100%. The placebo tablets were prepared by introducing 500 mg of the mixture. Measured responses were crushing strength (N), disintegration time (s), and price (cts / kg). A simplex lattice design for a special cubic model with three checkpoints was used. In this study, we consider both the variable dosage strength and the order of addition of components, focusing on a single response, i.e., tablet crushing strength (N). The OofA component-amount design for a quadratic model  (7) with 31 design points is used. The design is given in Table  7 with five levels of total amount i.e. 0, 250, 333.3, 375 and 500. The zero amount situation is used for the control test and is discussed in Piepel [15].
Table 7: The OofA mixture component-amount design for three ingredients and five levels of total amount AA.
Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA Runs $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA
@@
Now we use the TA algorithm and construct a G-optimal design with 18 optimal number of runs. We simulate the response variable (y), i.e., the crushing strength (N), for all design points in the TA algorithm-based G-optimal design. The design is given in Table 8. The average of the variances of predicted responses across all design points is 1, indicating that the overall uncertainty in the predictions is uniformly distributed throughout the design space. Since we have 89% G-efficiency, it implies that the design performs very well under the G-optimality criterion and ensures optimal precision even for the worst-case point in the design space. We fit the model  (7) to the design and use ANOVA to assess the significance of mixture amount and order effects on the response. The results are presented in Table  9. Using $\alpha$ = 0.05, it can be observed that component amounts and the order of addition have a significant effect on the crushing strength.
Table 8: G-optimal Component-amount OofA design obtained via TA algorithm
Run $a_{1}$ $a_{2}$ $a_{3}$ $z_{12}$ $z_{13}$ $z_{23}$ AA $y$
1 0 0 0 0 0 0 0 0
2 250 250 0 -1 0 0 500 63.5
3 250 250 0 1 0 0 500 120.4
4 0 500 0 0 0 0 500 98.5
5 250 0 250 0 -1 0 500 44.7
6 125 125 125 -1 1 1 375 68.7
7 500 0 0 0 0 0 500 77.2
8 250 0 250 0 1 0 500 130.6
9 0 250 0 0 0 0 250 46.3
10 0 250 250 0 0 1 500 68.6
11 0 0 250 0 0 0 250 50.4
12 125 125 125 1 1 -1 375 61.3
13 250 0 0 0 0 0 250 26.7
14 0 250 250 0 0 -1 500 49.0
15 125 125 125 -1 -1 -1 375 86.7
16 0 0 500 0 0 0 500 78.7
17 125 125 125 1 -1 -1 375 80.4
18 166.7 0 166.7 0 -1 0 333.3 39.4
Table 9: Parameter estimates for the OofA quadratic component–amount model
Term Estimate Std.Err t value Term Estimate Std.Err t value
$a_{0}$ 109.99 30.35 3.620 $a_{2}^{2}$ -4.20 15.77 -0.266
$a_{1}$ 56.65 27.57 2.054 $a_{3}^{2}$ -14.99 14.66 -1.022
$a_{2}$ 17.07 38.35 0.445 $a_{1}a_{2}$ 12.51 23.97 0.522
$a_{3}$ 61.02 29.57 2.063 $a_{1}a_{3}$ 5.54 23.41 0.236
$z_{12}$ 28.45 9.66 2.945 $a_{2}a_{3}$ -44.34 18.97 -2.337
$z_{13}$ 44.02 9.50 4.633 $a_{1}z_{12}$ 75.34 25.92 2.901
$z_{23}$ 9.80 9.66 1.014 $a_{3}z_{13}$ 92.86 25.93 3.582
$a_{1}^{2}$ 7.96 14.96 0.532 $a_{2}z_{23}$ 42.78 28.68 1.491
The FDS plot in Figure 3 shows that the design performs well across most of the space, with relatively low prediction variance for a large portion. The higher prediction variances are limited to a smaller portion of the design space.
Figure 3: FDS Plot for OofA component-amount model.
In a simulation study generating 10,000 designs, each with 36 runs, the Threshold Accepting (TA) algorithm was used to produce 16-run reduced designs. The full design had a mean G-efficiency value 1485.0, while the TA-reduced design achieved a mean G-efficiency value 934. The mean normalized G-efficiency for the TA design was 63%, indicating that the 16-run design retained approximately half of the predictive capacity of the full design. This demonstrates the TA algorithm’s effectiveness in selecting efficient design points, providing a substantial reduction in experimental runs without compromising performance, thereby validating its practical utility in experimental design.
Figure 4: Boxplot of MSEs across 10,000 simulated datasets
The Boxplot in Figure 4 shows that the TA-reduced design (16 points) achieves a lower median MSE than the full design (31 points), but with a wider spread of errors across simulations. This shows that the TA algorithm can substantially reduce the number of runs while maintaining predictive accuracy, although with increased variability due to the smaller sample size
@@ 5 Conclusion
In this study, we developed OofA mixture-amount experiments where the component amounts and the order of their addition to the mixture affect the overall response. We modified mixture-amount and component-amount models by incorporating the order-of-addition effect. To fit these models, we constructed designs for the OofA mixture-amount experiment. But such OofA mixture-amount designs have a large number of runs. To make the designs practical and cost effective, we used the TA algorithm that uses a subset of $t$ x $m!$ runs, which results in a much higher G-efficiency relative to the full OofA mixture-amount design. This reduction in run size improves design efficiency compared to the full OofA mixture-amount design, which contains many replicates. As recommended by [13], we have used the target run size $n\geq p$ , where $p$ is the number of model parameters.
The designs constructed in this paper can be used to determine whether the order and the mixture amounts have a significant effect on the response. These full mixture-amount OofA designs confirmed orthogonality between mixture-amount model terms and OofA effects. This allows for a precise estimation of both the parameters of the mixture components and the effects of order of addition.
The first example is the study of the joint action of three related hormones in groups of mice. An OofA mixture-amount design, with three levels of total amount A, was constructed. As the design has a large number of runs, we used the TA algorithm and selected a subset of runs, which is a G-optimal design with high G-efficiency value 89%. This indicates that the design is highly efficient and suitable for most practical applications.
In the second example, we generated a three-factor OofA component amount design with 31 runs and with five levels of total amount. We applied the TA algorithm and selected a partial fraction of the full design. It is a G-optimal design with the highest G-efficiency value 89% and a minimum number of runs. The design achieves a good trade-off between efficiency and robustness.
The developed OofA designs provide a systematic approach to include order-of-addition effects in mixture-amount experiments, enabling a more comprehensive design space exploration and better optimization of the response variable. It is the initial step for designing OofA mixture-amount experiments. We note that order-of-addition effects are most readily examined under controlled laboratory or pilot- scale conditions, and in practice, significant findings from such studies can inform the development of standardized procedures that facilitate consistent implementation during manufacturing. This work can be extended for constrained and orthogonally blocked mixture-amount experiments. Many other optimality criteria can be used for research interest. Metaheuristic algorithms, which do not require candidate points, can be used to construct optimal designs.
Acknowledgements
We thank the anonymous reviewers for their comments that helped us to improve the quality of the paper. Touqeer Ahmad acknowledges support from the Règion Bretagne through Project SAD-2021- MaEVa.
Conflict of interest
The authors declare no conflict of interest.
Authors contribution
Both authors equally contributed to this research.
Data availability statement
The data and code used in this study are available from the corresponding author upon reasonable request.
[…]